# GLDiTalker: Speech-Driven 3D Facial Animation with Graph Latent Diffusion Transformer

### v3

@@ 2.1 Speech-Driven 3D Facial Animation
Speech-driven talking face generation has attracted significant attention and achieved remarkable progress in recent years Karras et al. (2017); Richard et al. (2021); Fan et al. (2022); Xing et al. (2023); Peng et al. (2023a); Wu et al. (2023); Stan et al. (2023). Early methods, such as VOCA Cudeiro et al. (2019), introduce CNN-based frameworks for cross-identity face animation, while later works like Mimic Fu et al. (2024) and EmoTalk Peng et al. (2023b) focus on decoupling speech content, style, and emotion. Transformer-based approaches such as FaceFormer Fan et al. (2022) improve audio-visual mapping through attention mechanisms, and CodeTalker Xing et al. (2023) reduces cross-modal uncertainty using discrete motion priors. Recently, FaceDiffuser Stan et al. (2023) applies diffusion mechanisms to enhance the diversity of facial motion generation. Although these methods address various challenges, they still struggle to simultaneously achieve high accuracy and diversity due to limitations in modality alignment and inherent model constraints. To overcome these issues, we propose GLDiTalker, which introduces a quantized space-time diffusion training pipeline to better balance precision and diversity in speech-driven talking face generation.
@@ 2.2 Diffusion in Talking Head Animation
Inspired by the diffusion processes in non-equilibrium thermodynamics, Sohl-Dickstein et al. (2015) proposes diffusion, which differs from generative models like GAN Goodfellow et al. (2020), VAE Kingma and Welling (2014), and Flow Kingma and Dhariwal (2018). Diffusion is a Markov process capable of generating high-quality images from Gaussian noise, exhibiting a certain degree of stochasticity. Recently, many 2D talking head generation methods based on diffusion models have achieved highly realistic results Alghamdi et al. (2022); Guo et al. (2021); Liang et al. (2022); Park et al. (2022); Shen et al. (2022); Sun et al. (2022); Wang et al. (2023); Xu et al. (2024). For example, using audio as a condition, DREAM-Talk Zhang et al. (2023) designs a two-stage diffusion-based framework that generates diverse expressions while maintaining accurate audio-lip synchronization. DiffTalk Shen et al. (2023) further incorporates reference images and facial landmarks, enabling personality-aware synthesis and generalization across identities without fine-tuning. Similarly, VASA-1 Xu et al. (2024) employs diffusion in a latent space to model overall facial dynamics and head movements, effectively capturing subtle facial expressions and head motions. In the context of 3D facial animation, FaceDiffuser Stan et al. (2023) is the first to explicitly apply diffusion for speech-driven 3D facial motion generation. While diffusion methods inherently enhance diversity due to their stochastic nature, they often struggle to achieve high accuracy, particularly in terms of lip synchronization and motion precision. This trade-off between diversity and accuracy remains a significant challenge. To address this, we propose GLDiTalker, which leverages a quantized space-time diffusion training pipeline to balance diversity and accuracy.
@@ 3.1.1 Problem Definition
he facial motion $M_{1:T}=(m_{1},m_{2},...,m_{T})$ is a sequence of vertex displacements $m_{i}\in\mathbb{R}^{V\times 3}$ applied to a neutral template face mesh $f\in\mathbb{R}^{V\times 3}$, which consists of $V$ vertices. The audio $A_{1:T}=(a_{1},a_{2},...,a_{T})$ is a sequence of speech snippets, where each $a_{i}\in\mathbb{R}^{C}$ contains $C$ samples aligned with the corresponding motion $m_{i}$. Our objective is to predict a mesh sequence $\hat{M}_{1:T}$ based on the audio $A_{1:T}$ and the talking style represented by one-hot vectors $S=(s_{1},s_{2},...,s_{I})$, where $I$ denotes the number of speaking styles. The final face animations are obtained by summing the neutral template and the predicted motion, expressed as $\hat{F}_{1:T}={\hat{m}_{1+f},...,\hat{m}_{T+f}}$. The output sequence $\hat{F}_{1:T}$ is expected to achieve accurate lip synchronization with the input audio while exhibiting diverse and expressive facial motions that reflect different speaking styles.
@@ 3.1.2 Quantized Space-Time Diffusion Training Pipeline
To solve the problem of failing to simultaneously achieve high lip-sync accuracy and motion diversity caused by modality misalignment, GLDiTalker follows a two-stage pipeline that consists of a Graph Enhanced Quantized Space Learning Stage and a Space-Time Powered Latent Diffusion Stage, as illustrated in Fig. 2. In the first stage, GLDiTalker focuses on quantization to enhance lip-sync accuracy. Specifically, it employs a spatio-temporal VQ-VAE based on graph convolution and transformer to model facial motion into a quantized facial motion prior, ensuring precise alignment between audio and facial motion. In the second stage, the focus shifts to diffusion for generating diverse motions. A transformer-based denoising network is adopted for the reverse diffusion process, which iteratively transforms a standard Gaussian distribution into the facial motion prior. This process is conditioned on audio, speaker identity, and the diffusion step, enabling the generation of realistic and expressive facial animations. The animation results can be directly obtained from the denoising outputs by using the decoder from the first stage.
@@ 3.2 Stage 1: Graph Enhanced Quantized Space Learning for Lip-sync Accuracy Improvement
Existing methods have limited results on lip-sync, mostly limited by the ability to model the explicit space or the accumulation of errors due to autoregression. To improve the robustness to cross-modal uncertainty, GLDiTalker proposes a Graph Enhanced Quantized Space Learning Stage inspired by VQ-VAE Van Den Oord et al. (2017) to model the facial motion into a quantized motion prior, as illustrated in the first stage in Fig. 2. A graph based encoder is trained with both the temporal information dependency and the spatial connectivity of mesh vertices taken into account. Specifically, spatial and temporal information are serially processed to get latent feature $\hat{Z}\in\mathbb{R}^{T\times H\times C}$ with a novel designed Spatial Pyramidal SpiralConv Encoder $E^{V}$ following by a temporal transformer encoder $E^{T}$:
${E(\hat{M}_{1:T})}=E^{T}(E^{V}(\hat{M}_{1:T}))\rightarrow\hat{Z}_{1:T},$ (1)
@@ 4.5.2 Impact of Space-Time Powered Latent Diffusion Stage
The last column of the Tab. 3 shows that Graph Enhanced Quantized Space Learning Stage leads to significant improvements in terms of motion diversity (8.2176×$\times 10^{-4}$mm). With the same input, the deterministic model can only produce a fixed output, which does not reflect the many-to-many correspondence between speech and facial motions in reality. Instead, our approach uses diffusion in the quantized space, which even has an advantage of several orders of magnitude over FaceDiffuser, which adopts diffusion in the explicit space. We also conducted experiments to remove the Noise Encoder $E_{noise}$, which resulted in a performance degradation, as shown in Tab. 5, indicating the need of Noise Encoder.

### v4

@@ 2.1 Speech-Driven 3D Facial Animation
Speech-driven talking face generation has attracted significant attention and achieved remarkable progress in recent years Karras et al. (2017); Richard et al. (2021); Fan et al. (2022); Xing et al. (2023); Peng et al. (2023a); Wu et al. (2023); Stan et al. (2023); Shen et al. (2025b). Early methods, such as VOCA Cudeiro et al. (2019), introduce CNN-based frameworks for cross-identity face animation, while later works like Mimic Fu et al. (2024) and EmoTalk Peng et al. (2023b) focus on decoupling speech content, style, and emotion. Transformer-based approaches such as FaceFormer Fan et al. (2022) improve audio-visual mapping through attention mechanisms, and CodeTalker Xing et al. (2023) reduces cross-modal uncertainty using discrete motion priors. Recently, FaceDiffuser Stan et al. (2023) applies diffusion mechanisms to enhance the diversity of facial motion generation. Although these methods address various challenges, they still struggle to simultaneously achieve high accuracy and diversity due to limitations in modality alignment and inherent model constraints. To overcome these issues, we propose GLDiTalker, which introduces a quantized space-time diffusion training pipeline to better balance precision and diversity in speech-driven talking face generation.
@@ 2.2 Diffusion in Talking Head Animation
Inspired by the diffusion processes in non-equilibrium thermodynamics, Sohl-Dickstein et al. (2015) proposes diffusion, which differs from generative models like GAN Goodfellow et al. (2020), VAE Kingma and Welling (2014), and Flow Kingma and Dhariwal (2018). Diffusion is a Markov process capable of generating high-quality images from Gaussian noise, exhibiting a certain degree of stochasticity Shen and Tang (2024); Shen et al. (2025a). Recently, many 2D talking head generation methods based on diffusion models have achieved highly realistic results Park et al. (2022); Shen et al. (2022); Zhang et al. (2023); Shen et al. (2023); Xu et al. (2024). For example, using audio as a condition, DREAM-Talk Zhang et al. (2023) designs a two-stage diffusion-based framework that generates diverse expressions while maintaining accurate audio-lip synchronization. DiffTalk Shen et al. (2023) further incorporates reference images and facial landmarks, enabling personality-aware synthesis and generalization across identities without fine-tuning. Similarly, VASA-1 Xu et al. (2024) employs diffusion in a latent space to model overall facial dynamics and head movements, effectively capturing subtle facial expressions and head motions. In the context of 3D facial animation, FaceDiffuser Stan et al. (2023) is the first to explicitly apply diffusion for speech-driven 3D facial motion generation. While diffusion methods inherently enhance diversity due to their stochastic nature, they often struggle to achieve high accuracy, particularly in terms of lip synchronization and motion precision. This trade-off between diversity and accuracy remains a significant challenge. To address this, we propose GLDiTalker, which leverages a quantized space-time diffusion training pipeline to balance diversity and accuracy.
@@ 3.1.1 Problem Definition
The facial motion $M_{1:T}=(m_{1},m_{2},...,m_{T})$ is a sequence of vertex displacements $m_{i}\in\mathbb{R}^{V\times 3}$ applied to a neutral template face mesh $f\in\mathbb{R}^{V\times 3}$, which consists of $V$ vertices. The audio $A_{1:T}=(a_{1},a_{2},...,a_{T})$ is a sequence of speech snippets, where each $a_{i}\in\mathbb{R}^{C}$ contains $C$ samples aligned with the corresponding motion $m_{i}$. Our objective is to predict a mesh sequence $\hat{M}_{1:T}$ based on the audio $A_{1:T}$ and the talking style represented by one-hot vectors $S=(s_{1},s_{2},...,s_{I})$, where $I$ denotes the number of speaking styles. The final face animations are obtained by summing the neutral template and the predicted motion, expressed as $\hat{F}_{1:T}={\hat{m}_{1+f},...,\hat{m}_{T+f}}$. The output sequence $\hat{F}_{1:T}$ is expected to achieve accurate lip synchronization with the input audio while exhibiting diverse and expressive facial motions that reflect different speaking styles.
@@ 3.1.2 Quantized Space-Time Diffusion Training Pipeline
To solve the problem of failing to simultaneously achieve high lip-sync accuracy and motion diversity caused by modality misalignment, GLDiTalker follows a two-stage pipeline that consists of a Graph Enhanced Quantized Space Learning Stage and a Space-Time Powered Latent Diffusion Stage, as illustrated in Fig. 2. In the first stage, GLDiTalker focuses on quantization to enhance lip-sync accuracy. Specifically, it employs a spatio-temporal VQ-VAE Van Den Oord et al. (2017) based on graph convolution and transformer to model facial motion into a quantized facial motion prior, ensuring precise alignment between audio and facial motion. In the second stage, the focus shifts to diffusion for generating diverse motions. A transformer-based denoising network is adopted for the reverse diffusion process, which iteratively transforms a standard Gaussian distribution into the facial motion prior. This process is conditioned on audio, speaker identity, and the diffusion step, enabling the generation of realistic and expressive facial animations. The animation results can be directly obtained from the denoising outputs by using the decoder from the first stage.
@@ 3.2 Stage 1: Graph Enhanced Quantized Space Learning for Lip-sync Accuracy Improvement
Existing methods have limited results on lip-sync, mostly limited by the ability to model the explicit space or the accumulation of errors due to autoregression. To improve the robustness to cross-modal uncertainty, GLDiTalker proposes a Graph Enhanced Quantized Space Learning Stage inspired by VQ-VAE to model the facial motion into a quantized motion prior, as illustrated in the first stage in Fig. 2. A graph based encoder is trained with both the temporal information dependency and the spatial connectivity of mesh vertices taken into account. Specifically, spatial and temporal information are serially processed to get latent feature $\hat{Z}\in\mathbb{R}^{T\times H\times C}$ with a novel designed Spatial Pyramidal SpiralConv Encoder $E^{V}$ following by a temporal transformer encoder $E^{T}$:
${E(M_{1:T})}=E^{T}(E^{V}(M_{1:T}))\rightarrow\hat{Z}_{1:T},$ (1)
@@ 4.5.2 Impact of Space-Time Powered Latent Diffusion Stage
The last column of the Tab. 3 shows that using diffusion in the quantized space of Graph Enhanced Quantized Space Learning Stage leads to significant improvements in terms of motion diversity ($8.2176\times 10^{-4}$mm). With the same input, the deterministic model can only produce a fixed output, which does not reflect the many-to-many correspondence between speech and facial motions in reality. Instead, our approach uses diffusion in the quantized space, which even has an advantage of several orders of magnitude over FaceDiffuser, which adopts diffusion in the explicit space. We also conducted experiments to remove the Noise Encoder $E_{noise}$, which resulted in a performance degradation, as shown in Tab. 5, indicating the need of Noise Encoder.
@@ 5 Conclusion and Discussion
Acknowledgements
This work was supported by the National Natural Science Foundation of China under Grant No. 62441617. It was also supported by the Beijing Natural Science Foundation under Grant No. 4254100, the Fundamental Research Funds for the Central Universities under Grant No. KG16336301, the China Postdoctoral Science Foundation under Grant No. 2024M764093, and by Beijing Advanced Innovation Center for Future Blockchain and Privacy Computing.
Contribution Statement
Xiandong Li and Wenxiong Kang are the corresponding authors. They led the planning and execution of the research and provided financial support for the project leading to this publication. Yihong Lin and Zhaoxin Fan contributed equally to this work as co-first authors. Yihong Lin and Xiandong Li proposed the ideas. In addition, Yihong Lin designed the model and the implementation of the computer code, performed the experiments and wrote the initial draft. Zhaoxin Fan analyzed the study data and made critical review, commentary and revision including pre and post publication stages. Xianjia Wu and Lingyu Xiong assisted with figure preparation. Liang Peng, Songju Lei and Huang Xu proofread the manuscript. All authors have read and approved the final manuscript.